Please Help:
Find the future value of the annuity due.
$200 deposited at the beginning of each quarter for 7 years at 5% compounded quarterly
A) $6739.08
B) $6176.17
C) $1709.82
D) $6455.88

Answer by rfer(16322)

FV=(P*((1+i)^n-1)/i)(1+i)
FV=(200*((1+(.05/4)^28-1)/(.05/4))(1+i)
FV=((200*((1.0125)^28-1)/.0125)(1.0125)
FV=83.1984608/.0125=6655.876864*1.0125=6739.08 (A)